OpenStudy (anonymous):

To the nearest tenth, what is the area of the shaded segment when BN = 8 ft? A) 72.8 ft2 B) 39.3 ft2 C) 22.6 ft2 D) 53.2 ft2

OpenStudy (anonymous):

@tylermcmullen23

OpenStudy (anonymous):

\[\frac{ r^2 }{ 2 }(\frac{ \pi }{ 180 }c-\sin(c))\]

OpenStudy (anonymous):

thats the formula. so lets plug in the values

OpenStudy (anonymous):

do I put 8 in for r2?

OpenStudy (anonymous):

\[\frac{ 8^2 }{ 2 }(\frac{ \pi }{ 180 }*120-\sin(120))\]

OpenStudy (anonymous):

so yes

OpenStudy (anonymous):

i don't have a graphing calculator. but would it be 4(3.14/180 8120-sin(120))

OpenStudy (anonymous):

is it 7.5?

OpenStudy (anonymous):

without a graphing calculator, you woul do 8^2 which is 64 divided by 2 to get 32. so do 32(3.14/180*120-sin(120) yes

OpenStudy (anonymous):

but "/" meaning "divided by

OpenStudy (anonymous):

66.1? that's what I got but that's not an answer

OpenStudy (anonymous):

i got 39.3

OpenStudy (anonymous):

alright well thank you!!

OpenStudy (anonymous):

What is the area of the circle if XY = 10 ft? A) 20pi ft^2 B) 10pi ft^2 C) 5pi ft^2 D) 25pi ft^2

OpenStudy (anonymous):

anytime

OpenStudy (anonymous):

area is pi*r^2. it gives diameter, 10, si divide that by 2 to get 5

OpenStudy (anonymous):

5^2= ??? multiply that by pi

OpenStudy (anonymous):

multiply 25 by pi?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

78.5

OpenStudy (anonymous):

itd be d. i didnt know it was supposed to be in pi form lol

OpenStudy (anonymous):

haha okay so its 25pi?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Find the area of a square with a diagonal of 8 cm. A) 32 cm2 B) 18 cm2 C) 16 cm2 D) 12 cm2

OpenStudy (anonymous):

the formula is 1/2d^2

OpenStudy (anonymous):

so .5*8^2

OpenStudy (anonymous):

32

OpenStudy (anonymous):

What is the similarity ratio for two circles with areas 2pi m2 and 200pi m2? A) 1:10 B)1:100 C)1:2 D)1:5

OpenStudy (anonymous):

well, youd put 2:200. so simplify it. in this case, all you can do is divide by 2

OpenStudy (anonymous):

what am i dividing by 2?

OpenStudy (anonymous):

nvm its B, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Find the area of the trapezoid. A)307.5 in2 B) 365 in2 C) 615 in2 D)182.5 in2

OpenStudy (anonymous):

\[\frac{ a+b }{ 2 }h\]

OpenStudy (anonymous):

so \[\frac{ 20+29.2 }{ 2 }*12.5\]

OpenStudy (anonymous):

307.5?

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

Find the area of the trapezoid A) 307.5 in2 B) 365 in2 C) 615 in2 D)182.5 in2

OpenStudy (anonymous):

@tylermcmullen23

OpenStudy (anonymous):

we just done that one, remember? lol we got 307.5

OpenStudy (anonymous):

oops my bad

OpenStudy (anonymous):

its alright

OpenStudy (anonymous):

Both trapezoids are similar. The area of the smaller trapezoid is 13 ft2. Which is the best approximation for the area of the larger trapezoid? A) 20 ft2 B) 24 ft2 C) 28 ft2 D) 16 ft2

OpenStudy (anonymous):

first, you wanna find the scale factor from the small to the large. to do this, divide 10 by 8

OpenStudy (anonymous):

1.25

OpenStudy (anonymous):

since the area for the smaller is 13, multiply it by 1.25

OpenStudy (anonymous):

16.25

OpenStudy (anonymous):

so your answer should be 16! :)

OpenStudy (anonymous):

Both triangles are similar. The area of the smaller triangle is about 97 ft2. Which is the best approximation for the area of the larger triangle? 114 ft2 400 ft2 190 ft2 133 ft2

OpenStudy (anonymous):

135.8

OpenStudy (anonymous):

in this, i would say that 133 is your answer.. but im not entirely sure bc it didnt get as close as the trapezoid did.. thats weird

OpenStudy (anonymous):

yolo

OpenStudy (anonymous):

In circle P, PQ = 3 in. What is the circumference of the circle 36p in. 18p in. 6p in. 9p in.

OpenStudy (anonymous):

since circumference is found by doind diameter*pi.. multiply 3 by 2 to get the diameter, which is 6, and multiply it by 3.14

OpenStudy (anonymous):

18.84

OpenStudy (anonymous):

again, its supposed to be in pi form.. ._.

OpenStudy (anonymous):

so its just 18?

OpenStudy (anonymous):

so its 6p

OpenStudy (anonymous):

Thank you :-)

OpenStudy (anonymous):

One circular table has a diameter of 9 ft, and another circular table has a diameter of 14 ft. How much greater is the area of the larger table? Round to the nearest whole number. about 90 ft2 about 72 ft2 about 36 ft2 about 108 ft2

OpenStudy (anonymous):

well, the area for the smaller table is solved by doing 4.5^2*pi and for the larger, you solve by doing 7^2*pi

OpenStudy (anonymous):

153.86?

OpenStudy (anonymous):

yes that is correct for 7^2pi. but now, subtract 63.59from that

OpenStudy (anonymous):

90.27?

OpenStudy (anonymous):

Both trapezoids are similar. The area of the larger trapezoid is 61 ft2. Which is the best approximation for the area of the smaller trapezoid? 44 ft2 34 ft2 49 ft2 39 ft2

OpenStudy (anonymous):

id say d..

OpenStudy (anonymous):

Find the area of a regular hexagon whose apothem is 1.7 cm and side length is 2 cm. 3.4 cm2 10.2 cm2 20.4 cm2 6.8 cm2

OpenStudy (anonymous):

first, do the formula for a triangle using the given info. so do 2*1.7*1/2

OpenStudy (anonymous):

I got 1.7

OpenStudy (anonymous):

ok now how many triangles make up a hexagon??

OpenStudy (anonymous):

6?

OpenStudy (anonymous):

yes. now multiply that by the number you got, 1.7

OpenStudy (anonymous):

10.2

OpenStudy (anonymous):

To the nearest degree, what is the measure of the central angle for faucets? 43° 37° 48° 24°

OpenStudy (anonymous):

@tylermcmullen23

OpenStudy (anonymous):

im not entirely sure... but maybe 37. just because its a little under 45, closest to what i thought the answer was